OpenStudy (1645323):

use the unit circle to find the exact value of sin π/3

OpenStudy (anonymous):

root 3 over 2

OpenStudy (1645323):

Use the unit circle to find the exact value of tan 5pi/3. Thanks a lot btw

OpenStudy (1645323):

and sin pi/3

OpenStudy (anonymous):

negative root 3 over 2 divided by one-half = negative root 3

OpenStudy (1645323):

What is secθ if Sinθ=3/-5 and tanθ<0

OpenStudy (1645323):

is this too hard?

OpenStudy (anonymous):

sinx = y/r secx = r/y not sure if the tanx < 0 have to do with anything

OpenStudy (1645323):

the answer choices are only -2/5,2/5,-5/4,5/4

OpenStudy (1645323):

You are flying a kite on a string that is extended to 50 yards. If the kite is 23 yards in the air, what is the angle the kite is making with ground? Select one: a. 65 b. 38 c. 90 d. 52

OpenStudy (1645323):

what is arc sin? would the equation be sinx=23/50?